The marks of 500 candidates in an examination are normally distributed with a mean of 45


marks and a standard deviation of 20 marks.


(a) Given that the pass mark is 41, estimate the number of candidates who passed the


examination


(b) If 5% of the candidates obtain a distinction by scoring 𝑥 marks or more, estimate the value


of 𝑥.

1
Expert's answer
2022-01-14T06:50:13-0500

Let "Y=" the mark in an examination: "Y\sim N(\mu, \sigma^2)"

Given "\mu=45, \sigma=20."


a)



"P(Y\geq41)=1-P(Y<41)""=1-P(Z<\dfrac{41-45}{20})=1-P(Z<-0.2)""\approx0.579260""500(0.579260)=289.63"

289 candidates passed the examination.


b)



"P(Y\geq x)=1-P(Y<x)""=1-P(Z<\dfrac{x-45}{20})=0.05""P(Z<\dfrac{x-45}{20})=0.95""\dfrac{x-45}{20}=1.6449""x=78"